Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

Sales $ 80,000
Variable expenses 52,000
Contribution margin 28,000
Fixed expenses 21,840
Net operating income $ 6,160

6. If the selling price increases by $2 per unit and the sales volume decreases by 100 units, what would be the net operating income?

7. If the variable cost per unit increases by $1, spending on advertising increases by $1,700, and unit sales increase by 240 units, what would be the net operating income?

8. What is the break-even point in unit sales?

9. What is the break-even point in dollar sales?

10. How many units must be sold to achieve a target profit of $16,800?
Answer #1

Ans:

6. If the selling price increases by $2 per unit and the sales volume decreases by 100 units, what would be the net operating income?

S.NO

Particulars

Calculations

Amount($)

1

Sales

80,000/1000=80+2=82(1000-100)

73,800

2

Less: Variable cost

52000*900/1000

-46,800

3

Contribution Margin

27,000

4

Less:Fixed Expenses

-21,840

5

Net Operating Income

5,160

7. If the variable cost per unit increases by $1, spending on advertising increases by $1,700, and unit sales increase by 240 units, what would be the net operating income?

S.NO

Particulars

Calculations

Amount($)

1

Sales

80,000/1000=80(1000+240)

99,200

2

Less: Variable cost

52000/1000=52+1=53*1240

-65,720

3

Contribution Margin

33,480

4

Less:Fixed Expenses

21840+1700

-23,540

5

Net Operating Income

9,940

8. What is the break-even point in unit sales?

Break Even Point=Fixed costs/Selling price-variable cost

=21840/80-52=21840/28=780 Units

9. 9. What is the break-even point in dollar sales?

Break Even in value=Total Fixed cost/Contribution to sales ratio

=21840/.35=$62,400

Contribution to sales ratio=contribution/sales=28,000/80,000=35%

10. How many units must be sold to achieve a target profit of $16,800?

=Fixed Expenses + Target Profit/Contribution Margin per unit

=21840+16,800/28=1380 Units

Contribution Margin per unit=28000/1000=28